# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v2

@@ 7.2 Clustering on Incomplete Data
Clustering is an unsupervised learning task. Therefore, handling missing data in clustering is more challenging than in classification and regression. In recent decades, many clustering algorithms have been proposed. Typical clustering algorithms include k-means (ball1965isodata; IKOTUN2023178), DBSCAN (ester1996density), Gaussian mixture models, spectral clustering (shi2000normalized; ng2001spectral), subspace clustering (parsons2004subspace; elhamifar2013sparse), and deep learning-based clustering (xie2016unsupervised; ji2017deep; cai2022efficient; ren2024deep). When the data are incomplete, we can use some missing data imputation methods to fill the missing values and then perform these clustering algorithms. However, such a two-stage approach may not work well if the imputation algorithm does not align with the cluster structure in the data. There have been a few efforts in designing more effective approaches for clustering on incomplete data.
@@
sarkar2001fuzzy proposed an algorithm for fuzzy k-means on incomplete data, where the missing values are repaired incrementally in each iteration. wagstaff2004clustering presented a method for encoding partially observed features as a set of supplemental soft constraints and introduced an algorithm of k-means with soft constraints, which incorporates constraints into the clustering process, thereby eliminating the need for imputation. chi2016k developed an algorithm called k-POD for k-means with missing data. k-POD minimizes the sum of the squared differences between the data and the centers over the observed entries only, remaining unhindered by the need for accurate imputations. 8693952 proposed a variant of k-means that unifies clustering and imputation into one single objective, where missing values are optimized dynamically. The method outperformed k-means with kNN, EM, and mean imputations on a few datasets. armah2024k extended the idea to k-means with the Mahalanobis distance. There are also studies on kernel k-means for incomplete data such as (liu2020efficient) and Gaussian mixture models with missing data (zhang2021gaussian).
@@
Spectral clustering first constructs the similarity matrix and then performs k-means on the representations given by the eigenvalue decomposition of the Laplacian matrix (ng2001spectral). A simple strategy to handling missing data is computing the similarity between two data points using only the observed values, especially when the Gaussian kernel is used. However, this may not work well when the missing rate is high. lokse2017spectral developed a robust kernel function that is specifically designed to handle incomplete data and applied it to spectral clustering. 10586869 presented a spectral embedding fusion method for incomplete multi-view clustering. wen2022survey wrote a survey paper of incomplete multi-view clustering, which included a few extension methods of spectral clustering for incomplete data. NEURIPS2023_e5aa7171 proposed an imputation-free framework with two approaches to improving spectral clustering on incomplete data. The first approach is based on a kernel correction method that finds the best positive semi-definite approximation for the kernel matrix estimated on incomplete data. The second one is an extension of the first one, by using kernel self-expressive learning.
@@
Most subspace clustering methods can be organized into two categories. The first category is based on learning the subspaces directly (bradley2000k; fan2021large), while the second category is based on self-expressive models (elhamifar2013sparse) followed by spectral clustering. yang2015sparse developed two methods for subspace clustering on incomplete data. The first one generalizes the sparse subspace clustering algorithm so that it can obtain a sparse representation of the data using only the observed entries. The second one estimates a suitable kernel matrix by assuming a random model for the missing entries and obtains the sparse representation from this kernel. pimentel2016group also proposed two methods for subspace clustering on incomplete data: (a) group-sparse subspace clustering (GSSC), which is based on group-sparsity and alternating minimization, and (b) mixture subspace clustering (MSC), which models each data point as a convex combination of its projections onto all subspaces in the union. pimentel2016information analyzed the information-theoretic requirements of subspace clustering with missing data. fan2017sparse proposed to alternately compute the matrix of sparse representation coefficients and recover the missing entries of a data matrix. In particular, the algorithm recovers missing entries by minimizing the representation coefficients, representation errors, and matrix rank. tsakiris2018theoretical provided theoretical guarantees for sparse subspace clustering (elhamifar2013sparse) with incomplete data, showing that projecting the zero-filled data onto the observation pattern of the point being expressed can lead to substantial improvement in performance. lane2019classifying provided a comprehensive comparison of methods for subspace clustering with missing data. wang2020icmsc proposed a method for incomplete cross-modal subspace clustering. liu2021self developed an algorithm for self-representation subspace clustering on incomplete multi-view data. soni2025integer proposed to dynamically determine a set of candidate subspaces and optimally assign points to selected subspaces. The empirical study showed that the proposed approach can achieve high clustering accuracy even when the data are high rank, the percentage of missing data is high, or the subspaces are similar.
@@
de2019deep developed a variational deep embedding model for clustering of multivariate clinical patient trajectories with missing values. wen2020dimc proposed a deep learning model for incomplete multi-view clustering. xu2024deep provided a deep variational incomplete multi-view clustering algorithm that is able to explore shared clustering structures of incomplete data from different views. More work on deep clustering with missing data can be found in (ZHAO20181053; wen2021structural; xu2022deep; lin2022incomplete; tang2022deep).
@@ 7.3 Anomaly/Outlier/Novelty Detection on Incomplete Data
Anomaly detection (chandola2009anomaly), sometimes known as outlier detection (boukerche2020outlier), novelty detection (pimentel2014review), or even fault detection (in the manufacturing industry) (venkatasubramanian2003review), is usually an unsupervised learning task that aims to identify the samples deviating from common patterns. Typical algorithms include kernel density estimation, one-class SVM (scholkopf1999support), Isolation Forest (liu2008isolation), Deep SVDD (ruff2018deep), etc. Note that most of the dimensionality reduction methods (van2009dimensionality), such as PCA (abdi2010principal), KPCA (scholkopf1998nonlinear), and autoencoder (hinton2006reducing), can be used for anomaly detection or outlier detection. More papers about anomaly detection, especially those based on deep learning, can be found in (pang2021deep).
dietterich2018anomaly evaluated the detection performance of a few anomaly detection methods combined with different data imputation techniques. The experimental results showed that implementations contribute positively to the detection performance of unsupervised anomaly detection methods on incomplete data. fantii2022 investigated statistical process monitoring in the presence of missing data and introduced a fast incremental nonlinear matrix completion technique for online and sequential imputation. xu2021conformal developed ECAD, a method tailored for incomplete spatio-temporal data like traffic flow sequences, while ma2021probabilistic utilized Probabilistic Principal Component Analysis (PPCA) to recover missing values and identify anomalies. sarda2023unsupervised conducted a comparative analysis of existing unsupervised anomaly detection techniques applied to GAN-imputed datasets. These approaches follow a two-stage framework, in which the imputation models are trained exclusively on normal data or datasets with very few unlabeled outliers. Consequently, when applied to abnormal data during inference, the imputation model tends to reconstruct missing values based on learned normal patterns, thereby making anomalies appear more normal and reducing detection accuracy.
NEURIPS2024_f99f7b22 proposed an end-to-end method called ImAD for unsupervised anomaly detection in the presence of missing values. It integrates data imputation with anomaly detection into a unified optimization objective and automatically generates pseudo-abnormal samples to alleviate the imputation bias. They also proved the effectiveness of ImAD theoretically. zhang2025learning proposed an anomaly detection method for incomplete time series that flags unusual patterns by checking how consistently a two-step, patch-based imputation network fills the same series under different random masks.
@@ 7.4 Other Tasks on Incomplete Data
dimensionality reduction
(sanguinetti2006missing; nguyen2008robust; hu2010semiparametric)
(de2018nonlinear; gilbert2018unsupervised; ling2021dimension; yan2024inference)
change point detection
(xie2012change; londschien2021change; zhao2022online; follain2022high)
(liu2025high; xu2025online; enikeeva2025change)
causal inference & discovery
(kallus2018causal; yang2019causal; pmlr-v89-tu19a; nguyen2020bayesian)
(wang2020causal; huang2020causal; hillis2021causal)
(gao2022missdag; levis2025robust; landsiedel2025causal)
reinforcement learning
(yamaguchi2020model; mei2023reinforcement; chasalow2025missing)
@@ 8.1 Effective Parameter Estimation
dempster1977maximum proved that the EM algorithm for handling missing data converges to a local maximum point of the likelihood function. Although this does not directly guarantee accurate missing data imputation, it guarantees the effectiveness of the EM algorithm in estimating the parameters of many models such as mixture models and robust regression. chandrasekher2020imputation studied the problem of single imputation for high-dimensional sparse linear regression and obtained optimal consistency rates of both the LASSO (ranstam2018lasso) and the square-root LASSO without modification. bertsimas2024simple studied the performance of impute-then-regress pipelines by contrasting theoretical and empirical evidence. It established the asymptotic consistency of such pipelines for a broad family of imputation methods and showed that mean-impute is asymptotically optimal. verchand2024high presented an asymptotically exact characterization of the risk of ridge-regularized logistic regression when the observed covariates stem from a Gaussian error-in-variables model.
@@ 8.2 Exact Recovery Guarantee
CandesRecht2009 proved that a rank-$r$ matrix of size $n\times n$ can be recovered exactly via nuclear norm minimization when the number of observed entries (sampled uniformly at random) is larger than $Cn^{1.2}r\log n$ for some positive numerical constant $C$. The bound was further extended or improved by a few studies such as (candes2010matrix; MC_icml2014c1_chenc14). Regarding the low-rank factorization model for matrix completion, nonconvex but more efficient than nuclear norm minimization, a few scholars proposed alternating minimization and gradient descent algorithms that can exactly recover low-rank matrices (keshavan2012efficient; jain2013low; hardt2014understanding; ruoyu15). Compared to low-rank matrix completion, exact recovery guarantees for low-rank tensor completion is more difficult to obtain. Nevertheless, there have been a few guaranteed algorithms for tensor completion under CP decomposition (krishnamurthy2013low; mu2014square; jain2014provable; pmlr-v65-potechin17a; NEURIPS2020_dab1263d), Tucker decomposition (huang2014provable), and t-SVD (zhang2016exact).
@@ 8.3 Generalization Error Bound
srebro2005rank analyzed the generalization error bounds of rank, nuclear norm, and max-norm constrained matrix completion methods under a uniform sampling distribution. shamir14a provided a generalization bound for matrix completion with nuclear norm minimization under a milder assumption. (negahban2012restricted; fan2019factor) analyzed the recovery error bound for Schatten-$p$ quasi-norm based matrix completion. fan2021multi provided generalization bounds for deep matrix and tensor factorization models on incomplete data. fan2023euclideannorminduced analyzed the generalization bounds of Schatten-$p$ norm minimization for LRTC and showed that for LRTC with Schatten-$p$ quasi-norm regularizer on $d$-order tensors, $p=1/d$ is always better than any $p>1/d$ in terms of the generalization ability. wu2024smooth analyzed the generalization bound of total-variation regularized tensor completion. ledent2024generalization provided a tighter bound for matrix completion with Schatten-$p$ quasi-norm constraint.
@@ 9.1 Benchmarks and Toolboxes
Despite the extensive literature on missing data imputation, comprehensive toolboxes and benchmarks remain relatively limited. Several studies have sought to fill this gap. For instance, folch2016missing provided a MATLAB toolbox for missing data imputation and the imputation methods are mainly based on PCA and linear regression. josse2016missmda developed an R package for handling missing data. It includes principal component analysis for continuous variables, multiple correspondence analysis for categorical variables, factorial analysis on mixed data for both continuous and categorical variables, and multiple factor analysis for multi-table data. fancyimpute developed fancyimpute, a Python toolbox of missing data imputation, that included simple imputation, kNN, and a few LRMC methods. woznica2020does evaluated the empirical effectiveness of 7 imputation algorithms for 5 predictive models (logistic regression, classification tree, random forest, etc.) on 13 datasets. jager2021benchmark compared 6 imputation methods on 69 datasets under MCAR, MAR, and MNAR missingness patterns (with different missing rates) and evaluated by regression and classification downstream tasks. miao2022experimental experimentally compared 19 imputation algorithms on 15 real-world benchmark datasets. perez2022benchmarking compared 6 imputation methods in 10 classification and 3 regression tasks on 4 publicly available health databases. pereira2024imputation tested 6 imputation methods in an experimental setup that covers 10 datasets and 5 missingness levels (10% to 80%) in different MNAR settings. Empirically, they found that, for most missing rates and datasets, multiple imputation by chained equations performed best, whereas autoencoders showed promising results at higher missingness rates. cabrera2024benchmark provided a comparison of classical and deep learning imputation methods on 21 heterogeneous datasets in various areas. pons2024iti developed a toolbox to assess the reliability of various imputation methods, select the best imputer for any feature or group of features, and filter out features that do not meet quality criteria. toye2025benchmarking evaluated 12 imputation methods on two time-series datasets only. richter2025imputebench proposed ImputeBenc, which includes 4 imputation methods only, though more methods can be inserted into the workflow. There are more benchmarking papers on missing data imputation, most of which focus on specific data types or problems. For convenience, we summarize those published after 2020 in Table 6.
@@
(woznica2020does) general 7 13
(jager2021benchmark) general 6 69
(ahn2022comparison) time series 6 4
(miao2022experimental) general 19 15
(perez2022benchmarking) healthcare 6 4
(sun2023deep) general 6 10
kazijevs2023deep time series 7 5
(pereira2024imputation) general 6 10
(cabrera2024benchmark) general 8 21
(gama2024imputation) geological 8 3
(prakash2024benchmarking) health survey 4 10
(gendre2024benchmarking) biological 6 2
(du2024tsi) time series 28 8
(toye2025benchmarking) time series 12 2
(zhang2025comparison) healthcare 5 3
(richter2025imputebench) general 4

### v3

@@ 7.2 Clustering on Incomplete Data
Clustering is an unsupervised learning task. Therefore, handling missing data in clustering is more challenging than in classification and regression. In recent decades, many clustering algorithms have been proposed. Typical clustering algorithms include k-means (Ball and Hall, 1965; Ikotun et al., 2023), DBSCAN (Ester et al., 1996), Gaussian mixture models, spectral clustering (Shi and Malik, 2000; Ng et al., 2001), subspace clustering (Parsons et al., 2004; Elhamifar and Vidal, 2013), and deep learning-based clustering (Xie et al., 2016a; Ji et al., 2017; Cai et al., 2022b; Ren et al., 2024b). When the data are incomplete, we can use some missing data imputation methods to fill the missing values and then perform these clustering algorithms. However, such a two-stage approach may not work well if the imputation algorithm does not align with the cluster structure in the data. There have been a few efforts in designing more effective approaches for clustering on incomplete data.
@@
Sarkar and Leong (2001) proposed an algorithm for fuzzy k-means on incomplete data, where the missing values are repaired incrementally in each iteration. Wagstaff (2004) presented a method for encoding partially observed features as a set of supplemental soft constraints and introduced an algorithm of k-means with soft constraints, which incorporates constraints into the clustering process, thereby eliminating the need for imputation. Chi et al. (2016) developed an algorithm called k-POD for k-means with missing data. k-POD minimizes the sum of the squared differences between the data and the centers over the observed entries only, remaining unhindered by the need for accurate imputations. Wang et al. (2019) proposed a variant of k-means that unifies clustering and imputation into one single objective, where missing values are optimized dynamically. The method outperformed k-means with kNN, EM, and mean imputations on a few datasets. Armah and Melnykov (2024) extended the idea to k-means with the Mahalanobis distance. There are also studies on kernel k-means for incomplete data such as (Liu et al., 2020) and Gaussian mixture models with missing data (Zhang et al., 2021a).
@@
Spectral clustering first constructs the similarity matrix and then performs k-means on the representations given by the eigenvalue decomposition of the Laplacian matrix (Ng et al., 2001). A simple strategy for handling missing data is computing the similarity between two data points using only the observed values, especially when the Gaussian kernel is used. However, this may not work well when the missing rate is high. Løkse et al. (2017) developed a robust kernel function that is specifically designed to handle incomplete data and applied it to spectral clustering. Chen et al. (2024a) presented a spectral embedding fusion method for incomplete multi-view clustering. Wen et al. (2022) wrote a survey paper of incomplete multi-view clustering, which included a few extension methods of spectral clustering for incomplete data. Yu et al. (2023) proposed an imputation-free framework with two approaches to improving spectral clustering on incomplete data. The first approach is based on a kernel correction method that finds the best positive semi-definite approximation for the kernel matrix estimated on incomplete data. The second one is an extension of the first one, by using kernel self-expressive learning.
@@
Most subspace clustering methods can be organized into two categories. The first category is based on learning the subspaces directly (Bradley and Mangasarian, 2000; Fan, 2021), while the second category is based on self-expressive models (Elhamifar and Vidal, 2013) followed by spectral clustering. Yang et al. (2015) developed two methods for subspace clustering on incomplete data. The first one generalizes the sparse subspace clustering algorithm so that it can obtain a sparse representation of the data using only the observed entries. The second one estimates a suitable kernel matrix by assuming a random model for the missing entries and obtains the sparse representation from this kernel. Pimentel-Alarcón et al. (2016) also proposed two methods for subspace clustering on incomplete data: (a) group-sparse subspace clustering (GSSC), which is based on group-sparsity and alternating minimization, and (b) mixture subspace clustering (MSC), which models each data point as a convex combination of its projections onto all subspaces in the union. Pimentel-Alarcon and Nowak (2016) analyzed the information-theoretic requirements of subspace clustering with missing data. Fan and Chow (2017b) proposed to alternately compute the matrix of sparse representation coefficients and recover the missing entries of a data matrix. In particular, the algorithm recovers missing entries by minimizing the representation coefficients, representation errors, and matrix rank. Tsakiris and Vidal (2018) provided theoretical guarantees for sparse subspace clustering (Elhamifar and Vidal, 2013) with incomplete data, showing that projecting the zero-filled data onto the observation pattern of the point being expressed can lead to substantial improvement in performance. Lane et al. (2019) provided a comprehensive comparison of methods for subspace clustering with missing data. Wang et al. (2020a) proposed a method for incomplete cross-modal subspace clustering. Liu et al. (2021) developed an algorithm for self-representation subspace clustering on incomplete multi-view data. Soni et al. (2025) proposed to dynamically determine a set of candidate subspaces and optimally assign points to selected subspaces. The empirical study showed that the proposed approach can achieve high clustering accuracy even when the data are high rank, the percentage of missing data is high, or the subspaces are similar.
@@
de Jong et al. (2019) developed a variational deep embedding model for clustering of multivariate clinical patient trajectories with missing values. Wen et al. (2020) proposed a deep learning model for incomplete multi-view clustering. Xu et al. (2024b) provided a deep variational incomplete multi-view clustering algorithm that is able to explore shared clustering structures of incomplete data from different views. More work on deep clustering with missing data can be found in (Zhao et al., 2018; Wen et al., 2021; Xu et al., 2022; Lin et al., 2022; Tang and Liu, 2022).
@@ 7.3 Anomaly/Outlier/Novelty Detection on Incomplete Data
Anomaly detection (Chandola et al., 2009), sometimes known as outlier detection (Boukerche et al., 2020), novelty detection (Pimentel et al., 2014), or even fault detection (in the manufacturing industry) (Venkatasubramanian et al., 2003), is usually an unsupervised learning task that aims to identify the samples deviating from common patterns. Typical algorithms include kernel density estimation, one-class SVM (Schölkopf et al., 1999), Isolation Forest (Liu et al., 2008), Deep SVDD (Ruff et al., 2018), etc. Note that most of the dimensionality reduction methods (Van Der Maaten et al., 2009), such as PCA (Abdi and Williams, 2010), KPCA (Schölkopf et al., 1998), and autoencoder (Hinton and Salakhutdinov, 2006), can be used for anomaly detection or outlier detection. More papers about anomaly detection, especially those based on deep learning, can be found in (Pang et al., 2021).
Zemicheal and Dietterich (2019) evaluated the detection performance of a few anomaly detection methods combined with different data imputation techniques. The experimental results showed that implementations contribute positively to the detection performance of unsupervised anomaly detection methods on incomplete data. Fan et al. (2022) investigated statistical process monitoring in the presence of missing data and introduced a fast incremental nonlinear matrix completion technique for online and sequential imputation. Xu and Xie (2021) developed ECAD, a method tailored for incomplete spatio-temporal data like traffic flow sequences, while Ma et al. (2021) utilized Probabilistic Principal Component Analysis (PPCA) to recover missing values and identify anomalies. Sarda et al. (2023) conducted a comparative analysis of existing unsupervised anomaly detection techniques applied to GAN-imputed datasets. These approaches follow a two-stage framework, in which the imputation models are trained exclusively on normal data or datasets with very few unlabeled outliers. Consequently, when applied to abnormal data during inference, the imputation model tends to reconstruct missing values based on learned normal patterns, thereby making anomalies appear more normal and reducing detection accuracy.
Xiao and Fan (2024) proposed an end-to-end method called ImAD for unsupervised anomaly detection in the presence of missing values. It integrates data imputation with anomaly detection into a unified optimization objective and automatically generates pseudo-abnormal samples to alleviate the imputation bias. They also proved the effectiveness of ImAD theoretically. Zhang et al. (2025e) proposed an anomaly detection method for incomplete time series that flags unusual patterns by checking how consistently a two-step, patch-based imputation network fills the same series under different random masks.
@@ 7.4 Other Tasks on Incomplete Data
dimensionality reduction (Sanguinetti and Lawrence, 2006; Nguyen and Torre, 2008; Hu et al., 2010) (De Bodt et al., 2018; Gilbert and Sonthalia, 2018; Ling et al., 2021; Yan et al., 2024)
change point detection (Xie et al., 2012b; Londschien et al., 2021; Zhao et al., 2022; Follain et al., 2022) (Liu and Safikhani, 2025; Xu et al., 2025; Enikeeva and Klopp, 2025)
causal inference & discovery (Kallus et al., 2018; Yang et al., 2019; Tu et al., 2019; Nguyen and Gouno, 2020) (Wang et al., 2020b; Huang et al., 2020b; Hillis et al., 2021) (Gao et al., 2022; Levis et al., 2025; Landsiedel et al., 2025)
reinforcement learning (Yamaguchi et al., 2020; Mei et al., 2023; Chasalow et al., 2025)
@@ 8.1 Effective Parameter Estimation
Dempster et al. (1977) proved that the EM algorithm for handling missing data converges to a local maximum point of the likelihood function. Although this does not directly guarantee accurate missing data imputation, it guarantees the effectiveness of the EM algorithm in estimating the parameters of many models such as mixture models and robust regression. Chandrasekher et al. (2020) studied the problem of single imputation for high-dimensional sparse linear regression and obtained optimal consistency rates of both the LASSO (Ranstam and Cook, 2018) and the square-root LASSO without modification. Bertsimas et al. (2024) studied the performance of impute-then-regress pipelines by contrasting theoretical and empirical evidence. It established the asymptotic consistency of such pipelines for a broad family of imputation methods and showed that mean-impute is asymptotically optimal. Verchand and Montanari (2024) presented an asymptotically exact characterization of the risk of ridge-regularized logistic regression when the observed covariates stem from a Gaussian error-in-variables model.
@@ 8.2 Exact Recovery Guarantee
Candès and Recht (2009) proved that a rank-$r$ matrix of size $n\times n$ can be recovered exactly via nuclear norm minimization when the number of observed entries (sampled uniformly at random) is larger than $Cn^{1.2}r\log n$ for some positive numerical constant $C$. The bound was further extended or improved by a few studies such as (Candes and Plan, 2010; Chen et al., 2014). Regarding the low-rank factorization model for matrix completion, nonconvex but more efficient than nuclear norm minimization, a few scholars proposed alternating minimization and gradient descent algorithms that can exactly recover low-rank matrices (Keshavan, 2012; Jain et al., 2013; Hardt, 2014; Sun and Luo, 2015). Compared to low-rank matrix completion, exact recovery guarantees for low-rank tensor completion is more difficult to obtain. Nevertheless, there have been a few guaranteed algorithms for tensor completion under CP decomposition (Krishnamurthy and Singh, 2013; Mu et al., 2014; Jain and Oh, 2014; Potechin and Steurer, 2017; Liu and Moitra, 2020), Tucker decomposition (Huang et al., 2014), and t-SVD (Zhang and Aeron, 2016).
@@ 8.3 Generalization Error Bound
Srebro and Shraibman (2005) analyzed the generalization error bounds of rank, nuclear norm, and max-norm constrained matrix completion methods under a uniform sampling distribution. Shamir and Shalev-Shwartz (2014) provided a generalization bound for matrix completion with nuclear norm minimization under a milder assumption. (Negahban and Wainwright, 2012; Fan et al., 2019) analyzed the recovery error bound for Schatten-$p$ quasi-norm based matrix completion. Fan (2022b) provided generalization bounds for deep matrix and tensor factorization models on incomplete data. Fan et al. (2023) analyzed the generalization bounds of Schatten-$p$ norm minimization for LRTC and showed that for LRTC with Schatten-$p$ quasi-norm regularizer on $d$-order tensors, $p=1/d$ is always better than any $p>1/d$ in terms of the generalization ability. Wu and Fan (2024) analyzed the generalization bound of total-variation regularized tensor completion. Ledent and Alves (2024) provided a tighter bound for matrix completion with Schatten-$p$ quasi-norm constraint.
@@ 9.1 Benchmarks and Toolboxes
Despite the extensive literature on missing data imputation, comprehensive toolboxes and benchmarks remain relatively limited. Several studies have sought to fill this gap. For instance, Folch-Fortuny et al. (2016) provided a MATLAB toolbox for missing data imputation and the imputation methods are mainly based on PCA and linear regression. Josse and Husson (2016) developed an R package for handling missing data. It includes principal component analysis for continuous variables, multiple correspondence analysis for categorical variables, factorial analysis on mixed data for both continuous and categorical variables, and multiple factor analysis for multi-table data. Rubinsteyn and Feldman (2016) developed fancyimpute, a Python toolbox of missing data imputation, that included simple imputation, kNN, and a few LRMC methods. Woźnica and Biecek (2020) evaluated the empirical effectiveness of 7 imputation algorithms for 5 predictive models (logistic regression, classification tree, random forest, etc.) on 13 datasets. Jäger et al. (2021) compared 6 imputation methods on 69 datasets under MCAR, MAR, and MNAR missingness patterns (with different missing rates) and evaluated by regression and classification downstream tasks. Miao et al. (2022) experimentally compared 19 imputation algorithms on 15 real-world benchmark datasets. Perez-Lebel et al. (2022) compared 6 imputation methods in 10 classification and 3 regression tasks on 4 publicly available health databases. Pereira et al. (2024) tested 6 imputation methods in an experimental setup that covers 10 datasets and 5 missingness levels (10% to 80%) in different MNAR settings. Empirically, they found that, for most missing rates and datasets, multiple imputation by chained equations performed best, whereas autoencoders showed promising results at higher missingness rates. Cabrera-Sánchez et al. (2024) provided a comparison of classical and deep learning imputation methods on 21 heterogeneous datasets in various areas. Pons-Suñer et al. (2024) developed a toolbox to assess the reliability of various imputation methods, select the best imputer for any feature or group of features, and filter out features that do not meet quality criteria. Toye et al. (2025) evaluated 12 imputation methods on two time-series datasets only. Richter et al. (2025) proposed ImputeBenc, which includes 4 imputation methods only, though more methods can be inserted into the workflow. There are more benchmarking papers on missing data imputation, most of which focus on specific data types or problems. For convenience, we summarize those published after 2020 in Table 6.
In practice, a common and rigorous approach to benchmarking imputation methods involves the amputation of complete datasets. In this procedure, missing values are generated under controlled conditions, dictating the mechanism (MCAR, MAR, MNAR), the proportion of missing data, and the dependency structures among features. This simulation framework allows for objective performance comparison against the known ground truth.
The simulation-based benchmarking paradigm is supported by a relatively mature ecosystem of tools and studies. For instance, specialized libraries (e.g., pyampute (Schouten et al., 2018, 2022), missMethods (Rockel, 2022)) facilitate controlled data amputation and method implementation. Reviews and frameworks (e.g., MissMecha (Zhou et al., 2025)) provide structured guidance for generating and studying missing data under different mechanisms. Additionally, auxiliary tools for visualizing missingness patterns (e.g., missingno (Bilogur, 2018)) and testing assumptions (e.g., MissMech (Jamshidian et al., 2014) for MCAR testing) further strengthen the experimental pipeline.
@@
(Woźnica and Biecek, 2020) general 7 13
(Jäger et al., 2021) general 6 69
(Ahn et al., 2022) time series 6 4
(Miao et al., 2022) general 19 15
(Perez-Lebel et al., 2022) healthcare 6 4
(Sun et al., 2023) general 6 10
Kazijevs and Samad (2023) time series 7 5
(Pereira et al., 2024) general 6 10
(Cabrera-Sánchez et al., 2024) general 8 21
(Gama et al., 2024) geological 8 3
(Prakash et al., 2024) health survey 4 10
(Gendre et al., 2024) biological 6 2
(Du et al., 2024b) time series 28 8
(Toye et al., 2025) time series 12 2
(Zhang et al., 2025b) healthcare 5 3
(Richter et al., 2025) general 4